OpenStudy (anonymous):

When the center of Earth is 2*10^8 kilometers from the center of Mars, the force of gravity between the two planets is about 64.32*10^22 newtons. The mass of Earth is about 6*10^24 kilograms, and the mass of Mars is about 6.4*10^23 kilograms. Using these values, estimate the gravitational constant

OpenStudy (anonymous):

\[\color{green}{F = G\dfrac{m_1 m_2}{d^2} \\ 64.32 \times 10^{22}~\text{N} = G \dfrac{(6 \times 10^{24}~\text{kg} )(6.4 \times 10^{23})~\text{kg}}{(2 \times 10^8~\text{m})^2}}\] Solve for G.